OpenStudy (jamanorthi):

The end of a 1,000 foot kite string is anchored to the ground. the string represents a straight line.At a point 50 feet along the string from the anchor point, the string is 10 feet above the ground.how high is the kite flying? can some one help me plz

OpenStudy (the_fizicx99):

I would assume, im not sure. x/100 = 10/50

OpenStudy (the_fizicx99):

x/1000 *

OpenStudy (the_fizicx99):

typo D:

OpenStudy (the_fizicx99):

10/50 = x/1000 cross multiply, 10000/50 = 200 10 / 50 = 200 / 100 ?

OpenStudy (jamanorthi):

is it 200 feet?

OpenStudy (the_fizicx99):

I think its 200 :3

OpenStudy (jamanorthi):

okay thanks

OpenStudy (the_fizicx99):

yw